# Minimal Models for RNA Simulations

### v1

@@ Abstract
The increasing importance of RNA as a prime player in biology can hardly be overstated. It is suspected that the functions of RNA are linked to their structures and dynamics. Many of the problems in RNA, such as folding and RNA-RNA interactions that drive phase separation even in the absence of proteins, require cations. Because experiments alone cannot directly reveal the dynamics of cation-RNA interactions, well calibrated theory and computations are needed to predict how ions control the behavior of RNA. In this review, we outline the development of coarse-grained models at different resolutions with application to phase separation in low complexity sequences. We also describe folding of ribozymes and riboswitches with a focus on the impact of monovalent and divalent cations. We outline major challenges that need to be overcome to simulate complex problems such as assembly of ribosomes.
Department of Chemistry, The University of Texas at Austin, Austin, TX 78712, USA
School of Pharmacy, University of Nottingham, Nottingham, NG7 2RD, UK
Department of Chemistry, University at Buffalo, NY 14260, USA
@@ 1 Introduction
Although RNA sometimes referred to as the “dark matter", the spotlight on RNA biology has never been brighter with the award of Nobel prizes in medicine in two consecutive years, one in 2023 for the COVID-19 vaccine based on nucleoside base modifications of RNA 1, and another in 2024 for the role of microRNA in regulating gene expression 2; 3. These are the latest high profile examples in RNA biology which took off after the demonstration that RNA could function as enzymes over forty years ago 4; 5; 6. In the intervening years the functional repertoire of RNA has continuously expanded producing one surprise after another and with no end in sight. As is always the case in science, the discovery of new phenomena usually create a treasure trove of problems that need quantitative explanations, which in this field require both biophysical experiments 7; 8; 9 and computational approaches 10; 11; 12; 13; 14.
@@
The focus of this review is to highlight the utility of coarse-grained (CG) models in the folding of RNA and phase separation in low complexity sequences. Although not a substitute for atomically detailed simulations in water with explicit ions, there is a great need for CG models in understanding the folding of ribozymes (RNA enzymes) as well in the simulations of multiple chains that span a range of time and length scales. Over the last two decades several studies have shown that CG models, beginning with the initial reports on the mechanical force-induced unfolding of a RNA hairpin 10; 15, have been efficacious in making predictions that compare favorably with experiments. The CG models, at different resolutions, have been particularly important in elucidating the role of divalent cations in controlling RNA folding 16; 17; 18.
@@
For the purposes of this article, let us classify CG models into a few categories depending on the resolution of the RNA and the treatment of ions (Figure 1). (i) In the Single Site Interaction (SIS) model, each nucleotide is represented by a bead 19; 20. Electrostatic interactions associated with phosphate groups and the cations are modeled implicitly using the Debye-Hückel (DH) interactions 19; 20; 21. (ii) In the Three Interaction Site (TIS) model 10, each nucleotide has three sites representing phosphate, sugar, and base. Interactions involving charged groups (phosphate and metal ions) are treated explicitly 16 or implicitly 22, depending on the applications. (iii) If the monovalent (such as Na+ or K+) cation concentration exceeds the concentration of divalent cations greatly, as is the case in several folding experiments, the electrostatic interactions associated with monovalent cations could be treated using the DH potential where as divalent cations (Mg2+ or Ca2+) have to be explicitly treated for accurate description of RNA folding. The mixed description of cations when combined with the TIS model for RNA is referred to as the TIS-CAT model 23. (iv) There are models that use higher resolution (six or seven beads per nucleotide) in which charged interactions are either neglected or treated in an ad hoc manner. Most of these models were not designed to investigate RNA folding 24 or RNA-RNA interactions. Due to space limitations, we only discuss models that have made predictions for experiments or compared simulation results directly with experimental findings. Models that have been developed for the purpose of structure predictions25, while important, are beyond the scope of this review.
@@ 2 Single Interaction Site (SIS) Model
The Single Interaction Site (SIS) model represents each nucleotide by single interaction center 19. The SIS model for folded RNA retains chain connectivity and favorable attractive interactions between sites that stabilize the native fold. The initial formulation of the SIS model was primarily used to provide a theoretical framework to predict the outcomes of single molecule pulling experiments in which mechanical forces were to probe the folding landscapes of RNA constructs and ribozymes  26; 27; 28. We first describe applications to force-driven conformational changes in a riboswitch to demonstrate the versatility of the model and set the stage for phase separation studies.
@@
Response of Riboswitches to Mechanical Force: Riboswitches, consiting of an aptamer domain (AD) and an expression platform, contain RNA elements in the untranslated region of mRNA that allosterically regulate translation and transcription in bacteria by binding to metabolites with exquisite specificity 29; 30. Often transcription and riboswitch folding are coupled 31; 32; 33. The functional selectivity is best illustrated using the purine riboswitches as examples. For example, adenine (add) riboswitch 34 activates translation upon binding the metabolite purine, whereas the structurally similar pbuE adenine riboswitch 28 activates transcription in the presence of purine. Both of these are ON riboswitches, which means that translation or transcription is activated only upon binding of the metabolite. In contrast, OFF riboswitches shut down gene expression when the metabolite binds to the aptamer domain. The SAM-III riboswitch binds S-adenosylmethionine (SAM) and represses translation upon binding SAM which sequesters the Shine-Dalgarno ribosome binding sequence. Single molecule force spectroscopy using optical tweezers, performed either by applying constant force to the ends of the RNA or by increasing the force at a constant rate, have been used to assess the characteristics of the intermediates in riboswitches 35; 28. To complement these studies and to describe the structures of the intermediates that the folding landscape of riboswitches could be quantitatively mapped using SIS model 36; 37.
@@
In conventional single-molecular force spectroscopy, performed in the force ramp or force clamp mode, only states that are populated in the accessible force range can be characterized. Experiments performing force pulses during stretch-quench cycles could uncover the hidden states through which RNA folds. Let us start from a fully unfolded riboswitch, prepared at a high stretching force, $f_{s}$, and then quench the force to $f_{q}$ that is less than the critical force needed to unfold for a time period $t_{q}$ to initiate refolding. If $t_{q}$ is less than the folding time RNA would only access partially structured state. By repeating the process of stretching and quenching the force for varying time period ($t_{q}$) folding can be interrupted, and the states that are difficult to capture in a single or step wise quenching of force could be resolved. The SIS model simulations with the metabolite SAM bound showed that a spectrum of states was populated when $t_{q}$ was varied. Importantly, the simulations showed 36 that the refolding time as a function of force is accurately predicted using calculated free energy profile and Kramers rate theory 38; 39.
@@
Phase separation in repeat RNA sequences: More recently, the SIS model has been extended to probe phase separation in low complexity RNA repeat sequences, which cause several neurological and neuromuscular disorders, such as the Huntington disease 40. In vitro experiments 41 showed that the trinucleotide repeat sequences such as (CAG)n and (CUG)n undergo phase separation in which a dense phase coexists with a dilute phase. Importantly, phase separation occurs only if the number of repeats $n$ exceeds a critical value. To determine the mechanism of droplet formation using the SIS model, we adjusted the only parameter in the SIS model, corresponding to the energy scale for base-pair interactions, to reproduce the known structures of a short (CAG)2 duplex. Multichain simulations recapitulate the length and concentration dependence of the phase separation, in quantitative agreement with the in vitro experiments (Figure 2 a and b). The driving force for coacervate formation is the intermolecular base-pair interactions that has both an enthalpic and entropic contribution. Unexpectedly, it was found that once RNA molecules are recruited in the droplets, they undergo large conformational change, from a hairpin-like conformation in isolation to a stretched state, to form an extensive network of intermolecular interactions (Figure 2c). This soft network, in turn, constrains the RNA conformational fluctuation and mobility, which was rationalized using reptation dynamics 42. The SIS model has been adopted to simulate a number of repeat RNA sequences using the LAMMPS, which apparently enhances the computational efficiency 43. Like the previous work, simulations of a number of low complexity RNA sequences showed that the enhanced number of ways of forming Watson-Crick pairs render stability to the dense droplets.
@@
Based on the SIS model, a theory was advanced for anticipating the propensity of RNA to self-associate. It was discovered that the free energy gap separating the ground state and low lying excited states in repeat sequences are small 21. As a result, such sequences can readily unfold, thus exposing a stretch of nucleotides that can engage in inter molecular base pairing 44; 45, resulting in phase separation. In contrast, the free energy gap in heterogeneous sequences in which the 5 and 3 ends are in proximity are sufficiently great that the excited state has negligible population, which prevents unwarranted RNA-RNA interactions.
@@
The single bead representation of RNA has recently been used to investigate phase separations in mixture of RNA and proteins.20; 46; 47 In these models, interactions of RNA residues are parameterized to reproduce condensate formation with proteins, but the capability of such models to fold RNA into complex secondary and tertiary structures has not been demonstrated. Understanding the impact of RNA in influencing phase separation in proteins will also require a consistent model of RNA as well as interactions with ions.
@@
Figure 2: RNA phase separation at the molecular level using coarse-grained simulations. (a, b) The single interaction site model reproduces droplet formation (a) that occurs only when the RNA with $\sim$30 CAG repeats and RNA concentration exceeds critical concentration (b) 48. (c) Snapshots showing incorporation event of a (CAG)47 molecule (red) into a droplet. The dramatic conformational change that occurs can be seen in the accompanying secondary structure diagrams. Left: Initially, the isolated (CAG)47 chain formed a hairpin shape due to self-complementary base pairs. Middle: The RNA is unwound from one end to form new base pairs with other (CAG)47 chains present on the surface of droplet. Right: Once incorporated into the droplet, each RNA chain tends to elongate as formation of inter-chain base pairs dominate due to entropic gain.
@@ 3 Three Interaction Site (TIS) Model
In the Three Interaction Site (TIS) model each nucleotide consists of three distinct interaction sites – phosphate (P), sugar (S), and nucleobase (B) (Fig. 1). This allows the TIS model to capture the chemically distinct behavior of these moieties, which is essential for describing the sequence dependent structural dynamics. Since its first application in 2005 10, the TIS model has been widely used to study RNA folding in a variety of contexts. Examples include RNA folding induced by temperature jump 15 or mechanical forces 49, folding pathways shifting with changes in ion concentration on the folding pathways 50, and effects of molecular crowding51, and solvent viscosity 52. The three-bead representation has also been widely adopted in other RNA models 25; 53 and to investigate various aspects of DNA biophysics 54; 55.
@@
Over the years dramatic improvements have been within the TIS model by incorporating the nearest neighbor thermodynamics into the TIS representation 22. This was validated by quantitative comparison with experiments on RNA thermodynamics, such as melting of RNA hairpins and pseudoknots (PKs) (Figure 3a), which are essential building blocks of larger RNA molecules 22; 49. Recent work has focused on developing methods to accurately describe electrostatic interactions associated with monovalent and divalent cations and phosphate groups (Figure 1b). In most CG models of RNA, electrostatic interactions are modeled using the DH potential, which is sufficient for monovalent ions. In fact, for example, the TIS model with DH interactions accurately captured that folding pathways of a PK as a function of monovalent ion concentration, in accord with experiments 50. However, it was a long-standing limitation of CG models that the effects of divalent cations cannot be described accurately using the DH theory. This limitation was overcome in an important study that proposed a way to include both monovalent Na+, K+, Cl-) and divalent ions (Mg2+, Ca2+) in the TIS model explicitly 16. Simulations using this model captures the complicated ion-RNA electrostatic interactions, enabling visualization of how binding of individual cations facilitate self-assembly of RNA molecules.
@@
In a series of papers we illustrated the success of the TIS model with explicit ions by quantitatively reproducing experiments on the folding of the Azoarcus group I intron ribozyme 16; 56; 17. In particular, the simulations reproduced the Mg2+ concentration dependence of folding 16. Furthermore, specific Mg2+ binding sites agreed with the experimentally determined positions of Mg2+ 16; 56 (Figure 3b). This allowed the authors to confidently discuss how individual Mg2+ ions navigate the formation of secondary and tertiary structural motifs in order, which cannot be traced experimentally. The model also captured a subtle difference between Mg2+ and Ca2+. Ca2+ can drive RNA folding to some extent similar to Mg2+, but not enough to reach its enzymatically active conformation due to the slightly larger size of the ionic radius.
@@
It has been shown experimentally that RNA of a certain length with a complex fold often gets caught in kinetic traps and takes a long time to fold or misfold 57. The TIS simulation of the Azoarcus ribozyme has recently been extended to a kinetics study that showed how such multi-step folding and misfolding of the RNA occurs upon sudden changes in solution conditions, such as Mg2+ concentration jump 17. By simulating a series of Mg2+-induced folding events, it has been shown that the fate of the RNA molecules – whether they rapidly fold to the catalytically active state, get trapped in intermediate states, or even misfold – can be determined at a very early stage when the collapsing chain arrangement is driven by Mg2+ ion binding to specific sites (Figure 3c). Notably, the major misfolded state that appeared in the simulation was similar to misfolded structures of Tetrahymena ribozyme recently observed by cryo-electron microscopy 58, giving further confidence to the simulation model.
@@
In an important application of the TIS model, ions (Mg2+, K+, and F-) were included explicitly to investigate the Mg2+-dependent transition of a fluoride riboswitch 59 folding in the presence of K+ and F- ions 60. Kumar and Reddy investigated the folding of the aptamer domain (AD), which led them to propose a novel mechanism for the function of the fluoride riboswitch and F- encapsulation by the cationic pocket composed of three Mg2+ ions. Folding of the AD induces sequential binding of the two Mg2+ ions to the phosphate groups driven in part by a transition from an outer to an inner-shell coordination through dehydration (Fig 4a). The third Mg2+ and F- ion bind to the riboswitch in two steps. Binding of Mg2+ results in trigonal cationic pocket. Subsequently, Mg2+ and F- form a water-mediated ion pair (Mg-water-F).
@@
Let us highlight couple of additional applications of the TIS model. Simulations of the central domain of the bacterial ribosome, performed over a broad range of Mg2+ concentration, reproduced a number of experimental findings quantitatively (Figure 3d) 61. Strikingly, the TIS model simulations predicted not only the multi-pathway folding of G-quadruplex folding but also characterized the structures of the intermediates, thus complementing the experimental results 62. The handful of applications show that the combination of TIS model simulations and experiments could be used to fully determine the folding of large RNAs.
@@
Figure 3: Ion-driven thermodynamics and kinetics of RNA folding driven by ions. (a) Melting profile of BWYV pseudoknot measured in experiment (black)63 at 500 mM NaCl and TIS simulation (red) with implicit monovalent salt at the same concentration. (b) Nucleotide position dependent ion fingerprints in the folding of Azoarcus ribozyme, represented by local concentrations of Mg2+ (blue) and K+ (green). Comparisons with the Mg2+ binding sites resolved in X-ray crystal structure are shown. (c) Left: Kinetic partitioning mechanism in the folding of ribozyme. Differences in partial formations of secondary and tertiary structures at an early stage, driven by specific binding of Mg2+, lead to dramatic difference in the fate – folding to the native structure or misfolding. A major source of misfolding is caused by topological frustration where the position of J8/7 strand (red) relative to neighboring helices are opposite to that in in the native state. Right: The same mechanism of misfolding was recently observed in cryo-EM experiments of a similar ribozyme from Tetrahymena 58. The images were adopted from 17 under the CC-BY license. (d) Left: Conformational transitions driven by Mg2+ in the central domain of bacterial ribosome. Right: Comparison of the high density spots of Mg2+ around the central junction obtained in the simulation (grey) with the experimentally observed Mg2+ positions (red).

### v2

@@ Abstract
The increasing importance of RNA as a prime player in biology can hardly be overstated. Problems in RNA, such as folding and RNA-RNA interactions that drive phase separation, require cations. Because experiments alone cannot reveal the dynamics of cation-RNA interactions, well calibrated theory and computations are needed to predict how ions control the behavior of RNA. The perspective describes the development and use of coarse-grained models at different resolutions. We focus on single- and three-interaction site interaction models, in which electrostatic interactions are treated using a combination of explicit and implicit representations. Applications to the folding of ribozymes and riboswitches are discussed, with emphasis on the role of monovalent and divalent cations. We also discuss phase separation in low complexity sequences. Challenges in the simulation of complex problems such as ribosome assembly and RNA chaperones, requiring developments of models for RNA-protein interactions, are pointed out.
1. Department of Chemistry, The University of Texas at Austin, Austin, TX 78712, USA
2. School of Pharmacy, University of Nottingham, Nottingham, NG7 2RD, UK
3. Department of Chemistry, University at Buffalo, NY 14260, USA
@@ 1 Introduction
Although RNA sometimes referred to as the “dark matter" of biology, the spotlight on RNA biology has never been brighter. The award of Nobel prizes in medicine in two consecutive years, one in 2023 for the vaccines based on nucleoside base modifications of RNA 1, and another in 2024 for the role of microRNA in regulating gene expression 2; 3 highlight the importance of RNA. These are the latest high profile examples in RNA biology, which took off after the demonstration that RNA could function as enzymes over forty years ago 4; 5; 6. In the intervening years the functional repertoire of RNA has continues to expand producing one surprise after another and with no end in sight. As is always the case in science, the discovery of new phenomena create a treasure trove of problems that need quantitative explanations, which in the studies of RNA folding and RNA-RNA interactions require both biophysical experiments 7; 8; 9; 10 and computational approaches 11; 12; 13; 14; 15; 16; 17; 18; 19.
@@
Here, we focus exclusively on the utility of coarse-grained (CG) models in the folding of RNA and phase separation in low complexity sequences. Although not a substitute for atomically detailed simulations in water with explicit ions 20, there is also a need for CG models in understanding the folding of ribozymes (RNA enzymes) as well as in the simulations of multiple chains that span a range of time and length scales. hver the last two decades several studies have shown that CG models, beginning with the initial reports on the mechanical force-induced unfolding of a RNA hairpin 11; 21, have been efficacious in making predictions that compare favorably with experiments. The CG models, at different resolutions, have been particularly important in elucidating the role of divalent cations in controlling RNA folding 22; 23; 24; 19. It is also possible to convert a coarse-grained representation into atomic coordinates for further analysis 25.
@@
For the purposes of this article, let us classify CG models into a few categories depending on the resolution of the RNA and the treatment of ions (Figure 1). (i) In the Single Interaction Site (SIS) model, each nucleotide is represented by a single bead 26; 27; 28. Electrostatic interactions associated with phosphate groups and the cations are modeled implicitly using the Debye-Hückel (DH) interactions 26; 27; 29. (ii) In the Three Interaction Site (TIS) model 11, each nucleotide is represented using three sites corresponding to phosphate, sugar, and base. Interactions between charged groups (phosphate and metal ions) are treated explicitly 22 or implicitly 30, depending on the applications. (iii) Because the concentration of monovalent ions is higher than divalent ions under both physiological and in vitro experimental conditions, we developed a mixed description of the electrostatic interactions and combined it with the TIS model for RNA. In the resulting TIS-CAT 31 model, monovalent ions are treated using the DH potential whereas divalent cations (Mg2+ or Ca2+) are explicitly treated for the accurate description of RNA folding. (iv) There are models that use higher resolution (six or seven beads per nucleotide) in which stacking and base pairing interactions are treated in an ad hoc manner 18; 32; 33. In this review, our main focus is recent development of the first three categories. In addition, models that have been developed for the purpose of structure predictions34, while important, are beyond the scope of this review. We refer the reader to a more comprehensive review of various RNA simulations and earlier development of CG models 35.
@@ 2 Single Interaction Site (SIS) Model
The Single Interaction Site (SIS) model represents each nucleotide by single interaction center 26. The SIS model for folded RNA retains chain connectivity and favorable attractive interactions between sites that stabilize the native fold. The initial formulation of the SIS model provided a theoretical framework to predict the outcomes of single molecule pulling experiments in which mechanical forces probed the folding landscapes of RNA constructs and ribozymes  36.
@@
Response of Riboswitches to Mechanical Force: Riboswitches, consisting of an aptamer domain (AD) and an expression platform, contain RNA elements in the untranslated region of mRNA that allosterically regulate translation and transcription in bacteria by binding to metabolites with exquisite specificity 37; 38. Often transcription or translation and riboswitch folding are coupled 39; 40; 41; 42. The functional selectivity is best illustrated using the purine riboswitches as examples. For example, adenine (add) riboswitch 43 activates translation upon binding the metabolite purine, whereas the structurally similar pbuE adenine riboswitch 36 activates transcription in the presence of purine. Both of these are ON riboswitches, which means that translation or transcription is activated only upon binding of the metabolite. In contrast, OFF riboswitches shut down gene expression when the metabolite binds to the aptamer domain. The SAM-III riboswitch binds S-adenosylmethionine (SAM) and represses translation upon binding SAM which sequesters the Shine-Dalgarno ribosome binding sequence. Single molecule force spectroscopy using optical tweezers, performed either by applying constant force to the ends of the RNA or by increasing the force at a constant rate, were used to assess the characteristics of the intermediates in riboswitches 44; 36.
@@
In conventional single-molecular force spectroscopy, performed in either the force ramp or force clamp mode, only states that are populated in the accessible force range are characterized. Experiments performed using force pulses during stretch-quench cycles could uncover the hidden states through which RNA folds. Starting from a fully unfolded riboswitch, at a high stretching force, refolding can be initiated by quenching the force to a value less than the critical force needed to unfold for a time period $t_{q}$. If $t_{q}$ is less than the folding time, RNA would only access partially structured states. By repeating the process of stretching and quenching the force for different values of $t_{q}$ folding can be interrupted, and the states that are difficult to capture in a single or step wise quenching of force could be resolved. The SIS model simulations with the metabolite SAM showed that a spectrum of states was populated when $t_{q}$ was varied. Importantly, the simulations 45 showed that the refolding time as a function of force is accurately predicted using calculated free energy profile and Kramers rate theory 46.
@@
Phase separation in repeat RNA sequences: The SIS model has been extended to probe phase separation in low complexity RNA repeat sequences 47. In vitro experiments 10 showed that the trinucleotide repeat sequences such as (CAG)n and (CUG)n undergo phase separation in which a dense phase coexists with a dilute phase. Interestingly, phase separation occurs only if the number of repeats $n$ exceeds a critical value. These findings were explained using the SIS model, with a single unknown energy scale for base-pair interactions, that was adjusted to reproduce the known structures of a short (CAG)2 duplex 16. Multichain simulations recapitulate the length and concentration dependence of the phase separation, in quantitative agreement with the in vitro experiments (Figure 2 a and b). The driving force for coacervate formation is the intermolecular base-pair interactions that has both an enthalpic and entropic contribution. Unexpectedly, it was found that once RNA molecules are recruited in the droplets, they undergo large conformational change, from a hairpin-like conformation in isolation to a nearly stretched state, to form an extensive network of intermolecular interactions (Figure 2c). The formation of the soft network, in turn, constrains the RNA conformational fluctuation and mobility, which was rationalized using reptation dynamics 48. The SIS model has been adopted to simulate a number of repeat RNA sequences using LAMMPS, which reportedly enhances computational efficiency by speeding up the simulations by a factor of three to five 17. Like the previous work, simulations of a number of low complexity RNA sequences showed that the enhanced number of ways of forming Watson-Crick pairs render stability to the dense droplets.
@@
Based on the SIS model, a theory was advanced for predicting the propensity of RNA to self-associate. It was discovered that the free energy gap separating the ground state and low lying excited states in repeat sequences are small relative to $k_{B}T$ 29; 49. As a result, such sequences can readily unfold, thus exposing a stretch of nucleotides that can engage in inter molecular base pairing 50; 51, resulting in phase separation. In contrast, the free energy gap in heterogeneous sequences in which the 5 and 3 ends are in proximity are sufficiently large 29 that the excited state has negligible population, which prevents unwarranted RNA-RNA interactions.
@@
The single bead representation of RNA has recently been used to investigate phase separations in mixture of RNA and proteins.27; 52; 53 In these models, interactions of RNA residues are parameterized to reproduce condensate formation with proteins, but the capability of such models to fold RNA into complex secondary and tertiary structures has not been demonstrated. Understanding the impact of RNA in influencing phase separation in proteins will also require a consistent models of RNA as well as interactions with ions.
@@
Figure 2: RNA phase separation using coarse-grained simulations. (a, b) The single interaction site model reproduces droplet formation (a) that occurs only when the RNA with $\sim$30 CAG repeats and RNA concentration exceeds the critical concentration needed to cause phase separation (LLPS). (b) 16. (c) Snapshots showing incorporation event of a (CAG)47 molecule (red) into a droplet. Dramatic conformational changes occurs as can be seen in the accompanying secondary structure diagrams. Left: Initially, the isolated (CAG)47 chain formed a hairpin shape due to self-complementary base pairs. Middle: The RNA is unwound from one end to form new base pairs with other (CAG)47 chains present on the surface of droplet. Right: Once incorporated into the droplet, each RNA chain tends to elongate as formation of inter-chain base pairs dominate due to entropic gain.
@@ 3 Three Interaction Site (TIS) Model
In the Three Interaction Site (TIS) model, each nucleotide consists of three distinct interaction sites – phosphate (P), sugar (S), and nucleobase (B) (Fig. 1). This allows the TIS model to capture the chemically distinct behavior of these moieties. Since its first application in 2005 11, the TIS model has been widely used to study RNA folding in a variety of contexts. Examples include RNA folding induced by temperature jump 21 or mechanical forces 54, folding pathways shifting with changes in ion concentration on the folding pathways 55, and effects of molecular crowding56, and solvent viscosity 57. The three-bead representation has also been widely adopted in other RNA models 34; 58 and to investigate various aspects of DNA biophysics 59; 60.
@@
Over the years dramatic improvements have been within the TIS model by incorporating the nearest neighbor thermodynamics into the TIS representation 30. This was validated by quantitative comparison with experiments on RNA thermodynamics, such as melting of RNA hairpins and pseudoknots (PKs) (Figure 3a), which are essential building blocks of larger RNA molecules 30; 54. Recent work has focused on developing methods to accurately describe electrostatic interactions associated with monovalent and divalent cations and phosphate groups (Figure 1b). In most CG models, electrostatic interactions are modeled using the DH potential, which is sufficient for monovalent ions. For example, the TIS model with DH interactions accurately captured that folding pathways of a PK as a function of monovalent ion concentration, in accord with experiments 55. However, a limitation of such CG models is that the effects of divalent cations could not be described accurately using the DH theory. This limitation was overcome in an important study that proposed a way to include both monovalent (Na+, K+, Cl-) and divalent ions (Mg2+, Ca2+) in the TIS model explicitly 22. Simulations using this model capture the ion-RNA electrostatic interactions, enabling visualization of how binding of individual cations facilitate the self-assembly of RNA molecules 22; 61.
@@
In a series of papers, the success of the TIS model with explicit ions has been demonstrated by quantitatively reproducing experiments on the folding of the Azoarcus group I intron ribozyme 22; 62; 23.The simulations reproduced the Mg2+ concentration dependence of folding 22. Furthermore, specific Mg2+ binding sites agreed with the experimentally determined positions of Mg2+ 22; 62 (Figure 3b). Interestingly, the predictive power of various other methods on Mg2+ ion distributions on the surface of RNA molecules was evaluated at the recent CASP16, reflecting the importance and difficulty of the problem. In the TIS simulations, the trajectories were dissected to predict how individual Mg2+ ions facilitate the formation of secondary and tertiary structural motifs, which cannot be tracked experimentally. The simulations also showed that subtle difference between Mg2+ and Ca2+. Ca2+ can drive RNA folding but not enough to reach the enzymatically active conformation  22.
@@
Theoretical arguments and experiments have shown that RNA with a complex fold are kinetically trapped in native-like misfolded states, thus requiring considerable time to fold 63. The TIS kinetic simulations showed that folds Azoarcus ribozyme folds by a multi-step process involving a direct path to the native state as well as trapping in misfolded metastable native-like states upon a jump in the Mg2+ concentration 23. Analyses of a series of Mg2+-induced folding events, revealed that the fate of the RNA molecules – whether they rapidly fold to the catalytically active state, get trapped in intermediate states, or even misfold – occur at the very early stage coinciding with the chain collapse arrangement that is driven by Mg2+ ion binding to specific sites (Figure 3c). Notably, the major misfolded state that was identified in the simulation was similar to the misfolded structures of Tetrahymena ribozyme resolved by cryo-electron microscopy 64.
@@
In an important application of the TIS model, ions (Mg2+, K+, and F-) were included explicitly to investigate the Mg2+-dependent transition of a fluoride riboswitch folding in the presence of K+ and F- ions 65. Kumar and Reddy investigated the folding of the aptamer domain (AD), which led them to propose a novel mechanism for the function of the fluoride riboswitch and F- encapsulation by the cationic pocket composed of three Mg2+ ions65. Folding of the AD induces sequential binding of two Mg2+ ions to the phosphate groups driven in part by a transition from an outer to an inner-shell coordination through dehydration (Fig 4a). The third Mg2+ and F- ion bind to the riboswitch in two steps. Binding of the Mg2+ results in trigonal cationic pocket. Subsequently, Mg2+ and F- form a water-mediated ion pair (Mg-water-F).
@@
Let us highlight a couple of additional applications of the TIS model. Simulations of the central domain of the bacterial ribosome, performed over a broad range of Mg2+ concentrations, reproduced a number of experimental findings quantitatively (Figure 3d) 61. Strikingly, the TIS model simulations predicted not only the multi-pathway folding of G-quadruplex folding but also characterized the structures of the intermediates, thus complementing the experimental results 66. The handful of applications show that the combination of TIS model simulations and experiments could be used to fully determine the folding of large RNAs, which currently cannot be achieved by other methods.
@@
Figure 3: Ion-driven thermodynamics and kinetics of RNA folding. (a) Experimental melting profile of BWYV pseudoknot (black)67 at 500 mM NaCl and TIS simulation (red) with implicit monovalent salt at the same concentration. (b) Nucleotide position dependent ion fingerprints in the folding of Azoarcus ribozyme, represented by local concentrations of Mg2+ (blue) and K+ (green). Comparisons with the Mg2+ binding sites resolved in X-ray crystal structure are shown. (c) Left: Kinetic partitioning mechanism in the folding of ribozyme. Differences in partial formations of secondary and tertiary structures at an early stage, driven by specific binding of Mg2+, lead to dramatic difference in the fate – folding to the native structure or native metastable states. A major source of misfolding is due to topological frustration where the position of J8/7 strand (red) relative to neighboring helices are opposite to that in in the native state. Right: The same mechanism of misfolding was recently observed in cryo-EM experiments of a similar ribozyme from Tetrahymena 64. The images were adopted from 23 under the CC-BY license. (d) Left: Conformational transitions driven by Mg2+ in the central domain of bacterial ribosome. Right: Comparison of the high density spots of Mg2+ around the central junction obtained in the simulation (grey) with the experimentally observed Mg2+ positions (red).